OpenStudy (anonymous):

If 8k = 4, what is the value of k?

OpenStudy (anonymous):

k = 32 k = 4 k = 2 k = 0.5

OpenStudy (alexandervonhumboldt2):

divide both sides by 8: 8k/8=4/8 4/8=1/2 thus k=1/2

OpenStudy (anonymous):

Well, 0.5 is basically 1/2, so I am thinking that half of 8 is 4

OpenStudy (mynameisnemo):

To solve this, you need to get k alone. In this equation, k is being multiplied by 8 (8k). You need to cancel the 8 out by doing the opposite of multiplying 8 on both sides (Dividing). It would be k/8 and 4/8. The equation would be k= 0.5 or k= 1/2

OpenStudy (anonymous):

I agree with AlexandervonHumboldt2 and MyNameIsNemo